Fermi National Laboratory
 

Operation Summary for October 31, 2011 - November 2, 2011

 

Monday October 31

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE and NuMI experiments taking beam.

During the day shift, at 9 AM, operators reported a NuMI beam permit trip due to an indication of a spike in the gas flow through a beam detector; operators investigated and discovered the gas bottles were being changed. Operations established beam to the Meson FTBF experiment T-1015 at 10 AM. A MI person began a Recycler closed orbit study at 2:59 PM.

During the evening shift, the Recycler closed orbit study ended and operators rebooted the Recycler low level RF system and damper front end at 4:49 PM. Operations disabled beam to T-1015 at 10:09 PM.

Tuesday November 1

The midnight shift began with Operations monitoring beam to the NuMI and MiniBooNE experiments.

During the midnight shift, at 2:50 AM, the Linac Preaccelerator air temperature alarmed on a low temperature indication, but the dome temperature had not changed. Later, the Preaccelerator temperature jumped back to normal. At 7:41 AM, the Neutron Therapy Facility (NTF) began preparing to treat two patients later in the morning.

During the day shift, Operations established beam to T-1015 at 10:11 AM. Linac personnel accessed the H- Source dome (the backup source) at 11:23 AM. At 12:49 PM, a Linac person conducted an hour-long pulse study. Operations turned off beam to T-1015 at 3:12 PM for an access. Operations prepared MTA for access at 3:27 PM.

During the evening shift, Operations established beam to T-1015 at 4:30 PM. A Linac RF station (LRF4) tripped off for the first of three times in about 90 minutes; it reset. Operations disabled beam to T-1015 at 10 PM; this ended the experiment’s run for this year.

Wednesday November 2

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE experiments.

During the midnight shift, operators conducted a Booster injection tune at 2:36 AM. At 3:58 AM, the NuMI beam permit tripped off for the first of two times in 17 minutes when a torroid (E:TOR101) went above its trip limit; it reset.

Operations Coordinator (OpsCo) Plans for Today

The plans for today include Booster, Recycler, and MI g minus 2 (g-2) studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/02/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory